OpenStudy (anonymous):

Find the slope of the line joining (-6,7) and (-6,-4) but in the answers it says m=undefined , why is the slope undefined?

OpenStudy (espex):

A quick glance and I notice that the 'x' value is the same in both sets, this tells me that it is a vertical line.

OpenStudy (anonymous):

because youre gradient would be (-11/0). and anything divided by 0 is undefined...also this line would be parellel to the y-axis

hero (hero):

The slope isn't undefined. It's infinite

hero (hero):

Infinite means no definite number.

OpenStudy (anonymous):

argh! it is not, i repeat NOT infinite. it is simply not a number

hero (hero):

Shush Satellite

OpenStudy (anonymous):

\[\frac{1}{0}\] is not infinity. not even as a limit statement (which this is not) it is just not a number.

OpenStudy (anonymous):

@Hero...my calculator begs to differ man xD...it gives me a maths error...which for me is good enough to say that its undefined

OpenStudy (anonymous):

it is true that \[\lim_{x\rightarrow 0^+}\frac{1}{x}=\infty\] but \[\lim_{x\rightarrow 0^-}=-\infty\] so even as a limit statement it does not make any sense to say that you get infinity when dividing by zero

OpenStudy (anonymous):

it is like asking "how many miles to the gallon do you get on your bicycle" or "what is the price per ounce of free beer" the question makes no sense, and neither would the answer

OpenStudy (anonymous):

that link in utter nonsense. pure garbage. sorry, you cannot divide by zero, and you do not get "infinite" when you do so

hero (hero):

I divided by zero and met God

OpenStudy (chriss):

I'm going with Satellite on this one lol.